# An abstract framework for heterogeneous coupling: stability, approximation and applications

### v1

@@ Appendix A Construction and properties of the pseudo inverse
(B.1) $\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\|\widehat{\lambda}\|_{\Lambda}\lesssim\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|v\|_{V}}=\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|\Pi_{d}v\|_{V}}\frac{\|\Pi_{d}v\|_{V}}{\|v\|_{V}}\\ \leq\|\Pi_{d}\|_{V\to V}\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|\Pi_{d}v\|_{V}}=\|\Pi_{d}\|_{V\to V}\sup_{v\in V_{h}}\frac{b(v,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}v\|_{V}}.$
We also have
$b(w,\widehat{\lambda})=\langle B_{\delta}w,\widehat{\lambda}\rangle=\langle B_{\delta}B_{\delta}^{+}B_{\delta}w,\widehat{\lambda}\rangle=b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda}),$
@@
(B.2) $\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\|\Pi_{d}\|_{V\to V}\sup_{w\in V_{h}}\frac{b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
Now, for $z\in Z$ arbitrary, we can write
(B.3) $\sup_{w\in V_{h}}\frac{b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}=\sup_{w\in V_{h}}\frac{b(B_{\delta}^{+}B_{\delta}(w+z),\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
The idea is now to take $z$ such that $B_{\delta}^{+}B_{\delta}(w+z)\in\mathfrak{R}(\mathcal{K})$. To show that this is possible we observe that, by the definition of $\mathcal{K}$, given $\phi\in\Lambda_{\delta}^{\prime}$ we have that $\zeta=\phi-\mathcal{K}([\phi])\in\ker\mathcal{R}$. Then, for all $\lambda\in\Lambda_{\delta}$
$0=\langle\zeta,\Pi_{\sigma}\lambda\rangle=\langle\Pi_{\sigma}^{T}\zeta,\lambda\rangle,$
that is, $\zeta\in\ker\Pi_{\sigma}^{T}$. We have the following lemma.
Lemma B.1.
Letting $G_{\delta}:Z\to\Lambda_{\delta}^{\prime}$ denote the restriction of $B_{\delta}$ to $Z$, it holds that $\ker\Pi^{T}_{\sigma}=\mathfrak{R}(G_{\delta})$.
Proof.
If $\phi=G_{\delta}z=B_{\delta}z$ then $\Pi_{\sigma}^{T}\phi$ satisfies
$\langle\Pi_{\sigma}^{T}\phi,\lambda\rangle=\langle\phi,\Pi_{\sigma}\lambda\rangle=\langle B_{\delta}z,\Pi_{\sigma}\lambda\rangle=0$
and we then have $\mathfrak{R}(G_{\delta})\subseteq\ker\Pi_{\sigma}^{T}$. Let now $\phi\in\ker\Pi_{\sigma}^{T}$. On the one hand we can split $\phi$ as $\phi=G_{\delta}z+\phi^{\perp}$ with $\sigma^{*}(\phi^{\perp},G_{\delta}z^{\prime})=0$ for all $z^{\prime}\in Z$. On the other hand we have, for all $\lambda\in\Lambda_{\delta}$,
$0=\langle\Pi_{\sigma}^{T}\phi,\lambda\rangle=\langle\phi,\Pi_{\sigma}\lambda\rangle=\langle G_{\delta}z+\phi^{\perp},\Pi_{\sigma}\lambda\rangle=\langle\phi^{\perp},\Pi_{\sigma}\lambda\rangle.$
Then we have that ϕsuperscriptitalic-$\phi^{\perp}$ simultaneously verifies
$\sigma^{*}(\phi^{\perp},G_{\delta}z)=0,\ \forall z\in Z,\qquad\text{ and }\qquad\langle\phi^{\perp},\widehat{\lambda}\rangle=0,\ \forall\widehat{\lambda}\in\widehat{\Lambda}_{\delta}.$
We claim that this implies that $\phi^{\perp}=0$, which, in turn, implies that $\phi\in\mathfrak{R}(G_{\delta})$. To prove the claim we observe that
$\sigma^{*}(\phi^{\perp},G_{\delta}z)=\langle\Sigma^{-1}\phi^{\perp},G_{\delta}z\rangle=0,$
which implies $\mu=\Sigma^{-1}\phi^{\perp}\in\widehat{\Lambda}_{\delta}$. But we have that
$\sigma(\mu,\widehat{\lambda})=\langle\Sigma\Sigma^{-1}\phi^{\perp},\widehat{\lambda}\rangle=0,\ \forall\widehat{\lambda}\in\widehat{\Lambda}_{\delta}$
and this implies $\mu=0$ and, consequently $\phi^{\perp}=0$, which, in turn, implies $\phi\in\mathfrak{R}(G_{\delta})$. ∎
We conclude that $\ker\mathcal{R}=\mathfrak{R}(G_{\delta})$, so that for all $\phi$ we have, for some $z\in Z$, $\phi-\mathcal{K}[\phi]=G_{\delta}z$. Then we can choose $z^{*}\in Z$ as
$z^{*}=\mathop{\text{argmin}}_{z:\ G_{\delta}z=(\mathcal{K}-1)B_{\delta}w}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
We have, with ${\widetilde{\varphi}}=[B_{\delta}(w+z^{*})]$,
(B.4) $\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle=b(B_{\delta}^{+}B_{\delta}(w+z^{*}),\widehat{\lambda})=b(B_{\delta}^{+}B_{\delta}w+z^{*},\widehat{\lambda})=b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda}),$
as well as
(B.5) $\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})=m(\mathcal{K}[B_{\delta}(w+z^{*})],\mathcal{K}[B_{\delta}(w+z^{*}))=m(\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}),\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}))\\ =a(B_{\delta}^{+}\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}),B_{\delta}^{+}\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}))\lesssim\|B_{\delta}^{+}\Pi_{\sigma}^{T}B_{\delta}(w+z^{*})\|_{V}^{2}=\|B_{\delta}^{+}B_{\delta}(w+z^{*})\|^{2}_{V},$
@@
where we could write $\Pi_{\sigma}^{T}B_{\delta}(w+z^{*})=B_{\delta}(w+z^{*})$ thanks to our choice of $z^{*}$. Plugging (B.4) and (B.5) into (B.3), and the resulting bound into (B.2) we obtain
@@
$\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\|\Pi_{d}\|_{V\to V}\left(\sup_{w\in V_{h}}\min_{\begin{subarray}{c}z\in Z\\ G_{\delta}z=(\Pi_{\sigma}^{T}-1)B_{\delta}w\end{subarray}}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}\right)\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}.$
@@
$\sup_{w\in V_{h}}\min_{\begin{subarray}{c}z\in Z\\ G_{\delta}z=(\Pi_{\sigma}^{T}-1)B_{\delta}w\end{subarray}}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}}{\|B_{\delta}^{+}B_{\delta}w\|_{V}}.$
@@
$\|B_{\delta}^{+}B_{\delta}(w+z)\|_{V}\leq\|B_{\delta}^{+}B_{\delta}w\|_{V}+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}w\|_{V}\\ =\|B_{\delta}^{+}B_{\delta}w\|_{V}+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}B_{\delta}^{+}B_{\delta}w\|\\ \lesssim\left(1+\|B_{\delta}^{+}(\Pi_{\sigma}^{T}-1)B_{\delta}\|_{V\to V}\right)\|B_{\delta}^{+}B_{\delta}w\|_{V}.$
@@
$\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim C(V_{h},d,\sigma)\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}.$
Acknowledment
EB was partially funded by the EPSRC project EP/W007460/1, Integrated Simulation at the Exascale: coupling, synthesis and performance. SB is member of the Gruppo Nazionale di Calcolo Scientifico – Istituto Nazionale di Alta Matematica (GNCS–INdAM).

### v2

@@ Acknowledgement
(A.4) $\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\|\widehat{\lambda}\|_{\Lambda}\lesssim\sup_{w\in V_{\delta}}\frac{b(w,\widehat{\lambda})}{\|w\|_{V}}\leq\sqrt{K^{*}}\sup_{w\in V_{\delta}}\frac{b(w,\widehat{\lambda})}{\|\Pi_{d}w\|_{\widetilde{a}}}\frac{\|\Pi_{d}w\|_{\widetilde{a}}}{\|w\|_{\widetilde{a}}}\\ \leq\sqrt{K^{*}}\|\Pi_{d}\|_{\widetilde{a}\to\widetilde{a}}\sup_{w\in V_{\delta}}\frac{b(w,\widehat{\lambda})}{\|\Pi_{d}w\|_{\widetilde{a}}}=\sqrt{K^{*}}\|\Pi_{d}\|_{\widetilde{a}\to\widetilde{a}}\sup_{w\in V_{\delta}}\frac{b(w,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}}$,
where we used (4.9) and (A.1). We also have
[…]
@@
(A.5) $\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\sqrt{K^{*}}\|\Pi_{d}\|_{\widetilde{a}\to\widetilde{a}}\sup_{w\in V_{\delta}}\frac{b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}}$.
Now, for $z\in Z$ arbitrary, by the definition of $\widehat{\Lambda}_{\delta}$ we have that
$b(B_{\delta}^{+}B_{\delta}z,\widehat{\lambda})=b(z,\widehat{\lambda})=0.$
Then we can write
(A.6) $\sup_{w\in V_{\delta}}\frac{b(B_{\delta}^{+}B_{\delta}w,\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}}=\sup_{w\in V_{\delta}}\frac{b(B_{\delta}^{+}B_{\delta}(w+z),\widehat{\lambda})}{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{\widetilde{a}}}\frac{\|B_{\delta}^{+}B_{\delta}(w+z)\|_{\widetilde{a}}}{\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}}$.
The idea is now to take $z=z^{*}(w)$, with $z^{*}(w)$ such that $B_{\delta}(w+z^{*}(w))\in\mathfrak{R}(\mathcal{K})$. This is obtained by taking $B_{\delta}z^{*}=Gz^{*}=-Q_{\sigma}^{*}B_{\delta}w$. We have, with ${\widetilde{\varphi}}=[B_{\delta}w-Q_{\sigma}^{*}B_{\delta}w]=[B_{\delta}(w+z^{*}(w))]=[B_{\delta}w]$ (we recall that $\phi]$ stands for the equivalence class of $\phi$), and then
[…]
We now point out that for all $\phi\in\Lambda_{\delta}^{\prime}$ we have $\mathcal{K}([\phi])=\Pi_{\sigma}^{T}\phi$, which can be seen by comparing (4.4) and (4.6). We then can write
(A.8) $\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})=m(\mathcal{K}([B_{\delta}(w+z^{*}(w))]),\mathcal{K}([B_{\delta}(w+z^{*}(w))]))\\ =m(\Pi_{\sigma}^{T}(B_{\delta}(w+z^{*}(w))),\Pi_{\sigma}^{T}(B_{\delta}(w+z^{*}(w)))\\ =\widetilde{a}(B_{\delta}^{+}\Pi_{\sigma}^{T}(B_{\delta}(w+z^{*}(w))),B_{\delta}^{+}\Pi_{\sigma}^{T}(B_{\delta}(w+z^{*}(w))))\\ =\widetilde{a}(B_{\delta}^{+}B_{\delta}(w+z^{*}(w)),B_{\delta}^{+}B_{\delta}(w+z^{*}(w))),$
@@
where we could write $\Pi_{\sigma}^{T}B_{\delta}(w+z^{*}(w))=B_{\delta}(w+z^{*}(w))$ thanks to our choice of $z^{*}(w)$. Plugging (A.7) and (A.8) into (A.6), and the resulting bound into (A.5) we obtain
@@
$\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\sqrt{K^{*}}\|\Pi_{d}\|_{\widetilde{a}\to\widetilde{a}}\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}\sup_{w\in V_{\delta}}\ \frac{\|B_{\delta}^{+}B_{\delta}(w+z^{*}(w))\|_{\widetilde{a}}}{\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}}$.
@@
$\sup_{w\in V_{\delta}}\ \frac{\|B_{\delta}^{+}B_{\delta}(w+z^{*}(w))\|_{\widetilde{a}}}{\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}}=\sup_{w\in V_{\delta}}\frac{\|B_{\delta}^{+}(B_{\delta}w-Q_{\sigma}^{*}B_{\delta}w)\|_{\widetilde{a}}}{\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}}$.
@@
$\|B_{\delta}^{+}(B_{\delta}w-Q_{\sigma}^{*}B_{\delta}w)\|_{\widetilde{a}}\leq\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}+\|B_{\delta}^{+}Q_{\sigma}^{*}B_{\delta}w\|_{\widetilde{a}}\\ =\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}+\|B_{\delta}^{+}Q_{\sigma}^{*}B_{\delta}B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}\\ \lesssim\left(1+\|B_{\delta}^{+}Q_{\sigma}^{*}B_{\delta}\|_{\widetilde{a}\to\widetilde{a}}\right)\|B_{\delta}^{+}B_{\delta}w\|_{\widetilde{a}}$.
@@
$\sqrt{s_{h}(\widehat{\lambda},\widehat{\lambda})}\lesssim\sqrt{K^{*}}\|\Pi_{d}\|_{\widetilde{a}\to\widetilde{a}}\left(1+\|B_{\delta}^{+}Q_{\sigma}^{*}B_{\delta}\|_{\widetilde{a}\to\widetilde{a}}\right)\sup_{{\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}}\frac{\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle}{\sqrt{\widehat{m}({\widetilde{\varphi}},{\widetilde{\varphi}})}}$,
which concludes the proof.
Appendix B Proof of Lemma 5.5
We start by noting
$\|\phi\|_{s,\Gamma^{k}}^{2}=|\phi|^{2}_{s,\Gamma^{k}}+H^{1-2s}|\bar{\phi}|^{2}$
and the $H^{-s}$ norm as
$\|\phi\|^{2}_{-s,\Gamma^{k}}=|\phi|^{2}_{-s,\Gamma^{k}}+H^{2s-1}H^{2(d-1)}|\bar{\phi}|^{2}$
We have
(B.1) $|\phi|_{-1/2,\Gamma^{k}}=\sup_{\psi\in H^{1/2}(\Gamma^{k})}\frac{\sum_{F\subset\Gamma^{k}}\int_{F^{k}}(\phi\pm\bar{\phi}^{F})\psi}{|\psi|_{1/2,\Gamma^{k}}}\\ \leq\sup_{\psi\in H^{1/2}(\Gamma^{k})}\frac{\sum_{F\subset\Gamma^{k}}\int_{F^{k}}(\phi-\bar{\phi}^{F})\psi}{|\psi|_{1/2,\Gamma^{k}}}+\sup_{\psi\in H^{1/2}(\Gamma^{k})}\frac{\sum_{F\subset\Gamma^{k}}\int_{F^{k}}(\bar{\phi}^{F}-\bar{\phi})\bar{\psi}^{F}}{|\psi|_{1/2,\Gamma^{k}}}\\ \leq\sup_{\psi\in H^{1/2}(\Gamma^{k})}\frac{\sum_{F\subset\Gamma^{k}}|\phi|_{-1/2,F}|\psi|_{1/2,F}}{|\psi|_{1/2,\Gamma^{k}}}+\sup_{\psi\in H^{1/2}(\Gamma^{k})}\frac{\sum_{F\subset\Gamma^{k}}H^{d-1}|\bar{\phi}^{F}-\bar{\phi}||\bar{\psi}^{F}-\bar{\psi}|}{|\psi|_{1/2,\Gamma^{k}}}\\ \lesssim\sqrt{\sum_{F\subset\Gamma^{k}}|\phi|_{-1/2,F}^{2}}+\sqrt{\sum_{F\subset\Gamma^{k}}H^{2(d-1)}|\bar{\phi}^{F}-\bar{\phi}|^{2}}.$
This leads to
$|\phi|_{-1/2,\Gamma^{k}}^{2}\lesssim\sum_{F\subset\Gamma^{k}}|\phi|_{-1/2,F}^{2}+\sum_{F\subset\Gamma^{k}}H^{2(d-1)}|\bar{\phi}^{F}-\bar{\phi}|^{2}.$
Let us check the converse bound. There holds
$|\phi|_{-1/2,F}=\sup_{\psi\in H^{1/2},\fint\phi=0}\frac{\int_{F}\phi\psi}{|\psi|_{1/2,F}}\lesssim\sup_{\psi\in H^{1/2-\varepsilon},\fint\phi=0}\frac{\int_{F}\phi\psi}{|\psi|_{1/2-\varepsilon,F}}\\ \lesssim\underbrace{\frac{1}{\varepsilon}\sup_{\psi\in H^{1/2}}\frac{\int_{F}\phi\psi}{|\psi|_{H^{1/2-\varepsilon}_{0}(F)}}\lesssim\frac{1}{\varepsilon}\sup_{\psi\in H^{1/2}_{00}}\frac{\int_{F}\phi\psi}{|\psi|_{H^{1/2-\varepsilon}_{0}(F)}}}_{\text{density of $H^{1/2}_{00}$ in $H^{1/2-\varepsilon}_{0}$}}\lesssim\frac{h^{-\varepsilon}}{\varepsilon}\sup_{\psi\in H^{1/2}_{00}}\frac{\int_{F}\phi\psi}{|\psi|_{H^{1/2}_{00}(F)}}$
whence
$|\phi|_{-1/2,F}\lesssim\log(H/h)\|\phi\|_{H^{1/2}_{00}(F)^{\prime}}.$